# Argument Rarity-based Originality Assessment for AI-Assisted Writing

### v1

@@ 3.1 Framework Overview
This study proposes AROA, a structure-oriented originality assessment framework that evaluates essays based on the originality of achieved thinking rather than surface fluency or author attribution, in an environment where AI-assisted essay writing has become commonplace. Fig. 1 shows the overall processing pipeline of AROA.
@@
The design philosophy of AROA is based on Toulmin’s [31] argumentation structure model and integrates insights from critical thinking education [11, 12] and deep learning theory [5]. In particular, in the current situation where AI-generated text surpasses humans in surface fluency, the necessity of structurally evaluating learners’ cognitive depth is also supported from the perspective of knowledge construction theory [28].
@@
The rationale for selecting these four perspectives is based on the fact that originality is a multifaceted concept that cannot be captured by a single dimension. Structural rarity evaluates the uniqueness of how arguments are made, claim and evidence rarity evaluate the uniqueness of what is argued, and cognitive depth evaluates the uniqueness of how deeply arguments are developed. This multifaceted approach enables evaluation of originality accompanied by quality as argumentation, rather than superficial novelty. The following sections explain the details of each processing stage in order.
@@ 3.2 Argument Structure Extraction
This section explains the process of extracting argument structure from essays and converting it into a format suitable for subsequent rarity assessment. First, for each essay $x$, we define a structural representation consisting of the following five elements:
@@
The definition of each element is shown in Table 1. This five-element representation is an adaptation of Toulmin’s [31] argumentation structure model to the context of educational assessment. While Toulmin’s original model defines six elements, Claim, Data, Warrant, Backing, Qualifier, and Rebuttal, this study adopts a five-element model that integrates Warrant and Backing into Reasons and omits Qualifier. This simplification aims to improve practicality in educational settings and consistency of LLM extraction.
@@
These five elements form hierarchical support relationships. Reasons support Claims, and Evidences back up Reasons. Additionally, Counterarguments oppose Claims, and Rebuttals respond to Counterarguments. This relational structure enables capturing not only the presence or absence of elements but also the logical composition of argumentation.
An LLM is used for structure extraction. The input to the LLM consists of a two-layer structure: a system prompt and a user prompt. In the system prompt, the definition and guidelines for the extraction task were specified as follows:
@@
The three guidelines specified aim to ensure that extracted elements accurately reflect the intent of the original text, to unify element granularity by splitting compound sentences into semantic units, and to allow empty sets when elements do not exist. In the user prompt, the essay text was provided as input with instructions for structured output in JSON format. The output format includes text lists of each element as well as relational information between elements. To stabilize LLM output, the temperature parameter was set to $T=0$ to obtain deterministic output.
From the extracted structure, two derived features used in subsequent evaluation are calculated. First, the depth of reasoning $\text{Depth}(x)$ represents the hierarchical complexity of argumentation and is defined as follows:
@@
This definition progressively distinguishes from simple argumentation with only claims (depth 1) to complex argumentation with evidence backing or counterargument-rebuttal dialogic structure (depth 3).
Second, the number of counterargument-rebuttal matched pairs $n_{\text{matched}}$ is counted as the number of responds_to relationships extracted. When relationships are not explicitly extracted, it is estimated as $n_{\text{matched}}=\min(n_{A},n_{B})$. This feature captures the completeness of thinking, whether counterarguments are not only presented but also appropriately responded to.
@@
This study used a pre-trained Sentence-BERT model [14] to obtain $d=768$-dimensional embeddings. Sentence-BERT has been widely validated for its performance in capturing sentence-level semantic similarity and is suitable for this task. These embedding representations are used as the computational basis for semantic similarity in the claim and evidence rarity assessments explained in subsequent sections.
@@ 3.3 Structural Rarity Assessment
This section explains the method for calculating the structural rarity score $S_{\text{struct}}$, which quantifies how rare an essay’s argument structure is within the corpus. First, a 12-dimensional structural feature vector $\boldsymbol{\phi}(x)$ is constructed from each essay $x$:
@@
Here, $n_{C},n_{R},n_{E},n_{A},n_{B}$ are the counts of each element, $n_{R}/\max(n_{C},1)$ is the number of reasons per claim, and $n_{E}/\max(n_{R},1)$ is the number of evidences per reason. $\text{Depth}(x)$ and $n_{\text{matched}}$ are the depth of reasoning and the number of counterargument-rebuttal matched pairs defined in the previous section. $\text{Branches}(x)=\max(n_{C},1)\times\max(n_{R},1)$ represents the number of argument branches, capturing multifaceted argument structures that support multiple claims with multiple reasons. $\mathbb{I}_{n_{A}>0}$ and $\mathbb{I}_{n_{B}>0}$ are indicator functions showing the presence or absence of counterarguments and rebuttals. These features are designed to capture multiple aspects: quantitative aspects of argumentation (element counts), structural aspects (ratios, depth, branches), and critical thinking aspects (existence and correspondence of counterargument-rebuttal).
The structural rarity score is defined as the negative logarithm of probability density in the feature vector space. This definition corresponds to self-information in information theory, where essays located in low-density regions receive higher scores. Prior to density estimation, to eliminate scale differences between features, standardization is performed on the set of feature vectors across the entire corpus $\Phi=\{\boldsymbol{\phi}(x_{i})\}_{i=1}^{N}$:
@@
Here, $\boldsymbol{\mu}_{\Phi}$ and $\boldsymbol{\sigma}_{\Phi}$ are the corpus-wide mean and standard deviation of each feature. Next, to mitigate the curse of dimensionality associated with density estimation in high-dimensional spaces, dimensionality is reduced to 3 using Principal Component Analysis (PCA). Kernel Density Estimation (KDE) is used for density estimation, with bandwidth automatically set based on Scott’s rule. The final structural rarity score is defined as follows:
@@
Here, $\hat{p}(\boldsymbol{\phi}^{\prime}(x))$ is the KDE density estimate for the feature vector after standardization and PCA. This score causes essays with typical argument structures within the corpus to receive low scores, while essays with rare argument structures receive high scores.
@@ 3.4 Semantic Rarity Assessment
This section explains the method for calculating the semantic rarity scores $S_{\text{claim}}$ and $S_{\text{evid}}$, which quantify how rare the content of claims and evidence are within the corpus. While the structural rarity in the previous section evaluates how arguments are made, the semantic rarity in this section evaluates what is argued.
@@ 3.4.2 Evidence Rarity
These semantic rarity scores result in low scores for essays that use typical claims or evidence that frequently appear within the corpus, and high scores for essays that use original perspectives or rare evidence.
@@ 3.5 Cognitive Depth Assessment
This section explains the method for calculating the cognitive depth score $S_{\text{move}}$, which evaluates the depth of reasoning and completeness of response to counterarguments in argumentation. While the rarity assessments in previous sections measure whether something is rare, the cognitive depth in this section evaluates how deeply arguments are developed.
@@
The first component, $n_{\text{matched}}$, is the number of pairs where counterarguments and rebuttals semantically correspond. As explained in Section 3.2, this value is extracted as responds_to relationships during structure extraction, or estimated as $n_{\text{matched}}=\min(n_{A},n_{B})$ when relationships are not explicitly extracted. Therefore, a higher number of counterargument-rebuttal pairs indicates more active back-and-forth of critical thinking.
The second component, $\text{Depth}(x)$, is the depth of reasoning defined in Section 3.2, with 1 for claims only, 2 when reasons exist, and 3 when evidence exists or both counterarguments and rebuttals exist. This component evaluates hierarchical argument structures where claims are supported by reasons and further backed by evidence.
The third component, $\text{Coverage}(x)$, is the counterargument coverage rate, representing the proportion of presented counterarguments that are responded to with rebuttals:
@@
This metric captures the completeness of thinking, i.e., whether counterarguments are not only raised but also appropriately responded to. Essays that provide rebuttals to raised counterarguments rather than leaving them unaddressed receive higher scores.
This study used $\alpha=\beta=\gamma=1.0$. These weights represent a design that treats the three components equally, but they can be adjusted according to educational objectives. For example, if critical thinking is emphasized, $\alpha$ and $\gamma$ can be increased, and if hierarchical argumentation is emphasized, $\beta$ can be increased.
@@ 3.6 Quality Adjustment
This section explains the method for calculating the quality score $Q(x)$, which prevents essays that are original but logically flawed from receiving high scores. Since essays that are logically incoherent cannot be called original even if they have high rarity, adjustment by quality is necessary. The quality score consists of two components: Coherence and Logical Validity:
@@
As an important design decision, this framework intentionally excludes Fluency from quality evaluation. The reason is that text generated by LLMs generally has high fluency, so including fluency in the evaluation would favor AI-generated text, contradicting the objective of this framework to evaluate originality of thinking. Each component is evaluated using an LLM. For coherence evaluation, judgment is made based on three criteria: i) whether there are contradictions between claims and reasons, ii) whether there are contradictions between reasons and evidence, and iii) whether there are contradictions among multiple claims. Output is in three levels: 1.0 (no contradictions), 0.5 (partial contradictions), and 0.0 (serious contradictions). For logical validity evaluation, judgment is made based on three criteria: i) whether reasons logically support claims, ii) whether evidence appropriately backs up reasons, and iii) whether there are logical leaps or fallacies. Output is in three levels: 1.0 (valid), 0.5 (partial problems), and 0.0 (serious defects).
In both evaluations, the prompt explicitly instructs not to evaluate fluency or style, ensuring that evaluation focuses only on logical structure. This evaluates only the quality as argumentation rather than the superficial smoothness of text.
@@ 3.7 Final Score Computation
This section explains the method for integrating the four originality sub-scores and quality score to compute the final evaluation value $S_{\text{final}}(x)$. The four sub-scores ($S_{\text{struct}}$, $S_{\text{claim}}$, $S_{\text{evid}}$, $S_{\text{move}}$) each have different scales, making direct comparison and integration difficult. To solve this problem, each sub-score is z-score normalized within the corpus:
@@
Here, $\mu_{\text{comp}}$ and $\sigma_{\text{comp}}$ are the corpus-wide mean and standard deviation of each component. This normalization transforms each sub-score to a distribution with mean 0 and standard deviation 1, enabling fair comparison and integration between components.
@@
Here, $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ are the weights for each component, satisfying $\sum_{i=1}^{4}\lambda_{i}=1$. This study used $\lambda_{1}=\lambda_{2}=\lambda_{3}=\lambda_{4}=0.25$, a design that treats the four components equally.
@@
The multiplicative design ensures that when $Q(x)\approx 0$, $S_{\text{final}}(x)\approx 0$ even if originality is high, preventing logically flawed essays from receiving high scores. On the other hand, when $Q(x)\approx 1$, the originality score is almost directly reflected in the final score. This multiplicative design positions quality as a prerequisite for originality.
Finally, Algorithm 1 shows the overall processing procedure of AROA. Through this series of processes, AROA enables evaluation of essays based on the originality of achieved thinking rather than surface fluency or author attribution, in an environment where AI-assisted essay writing has become commonplace.
@@
17:  $\text{Coverage}(x_{i})\leftarrow n_{\text{matched}}(x_{i})/\max(n_{A}(x_{i}),1)$
18:  $S_{\text{move}}(x_{i})\leftarrow\alpha\cdot n_{\text{matched}}(x_{i})+\beta\cdot\text{Depth}(x_{i})+\gamma\cdot\text{Coverage}(x_{i})$
19:end for
20:// Stage 4: Quality Adjustment
21:for each $x_{i}\in\mathcal{D}$ do
22:  $Q_{\text{coherence}}(x_{i})\leftarrow\text{LLM\_Evaluate}(x_{i},\text{``coherence''})$
23:  $Q_{\text{logical}}(x_{i})\leftarrow\text{LLM\_Evaluate}(x_{i},\text{``logical''})$
24:  $Q(x_{i})\leftarrow w_{\text{coh}}\cdot Q_{\text{coherence}}(x_{i})+w_{\text{log}}\cdot Q_{\text{logical}}(x_{i})$
25:end for
26:// Stage 5: Final Score Computation
27:for each component $\text{comp}\in\{\text{struct},\text{claim},\text{evid},\text{move}\}$ do
28:  $\mu_{\text{comp}},\sigma_{\text{comp}}\leftarrow\text{Mean}(\{S_{\text{comp}}(x_{i})\}),\text{Std}(\{S_{\text{comp}}(x_{i})\})$
29:  $\tilde{S}_{\text{comp}}(x_{i})\leftarrow(S_{\text{comp}}(x_{i})-\mu_{\text{comp}})/\sigma_{\text{comp}}$ for all $i$
30:end for
31:for each $x_{i}\in\mathcal{D}$ do
32:  $\tilde{S}_{\text{orig}}(x_{i})\leftarrow\sum_{j=1}^{4}\lambda_{j}\tilde{S}_{j}(x_{i})$
33:  $S_{\text{final}}(x_{i})\leftarrow\tilde{S}_{\text{orig}}(x_{i})\times Q(x_{i})$
@@
35:return $\{S_{\text{final}}(x_{i})\}_{i=1}^{N}$
@@ 4.1 Dataset
This study used the AI Detection for Essays Dataset (AIDE) [6]. AIDE is a dataset constructed for AI-generated text detection in educational settings and contains both argumentative essays written by students and essays generated by AI. The essays concern two topics: Car-free cities and Electoral College, both of which are assignments requiring students to read given materials and argue their own position. There are two reasons for selecting this dataset. First, it contains numerous human essays on the same topic and same assignment, making it suitable for evaluating the proposed framework based on rarity within a corpus. Second, since it contains both human essays and AI essays, it enables verification of whether the proposed method can distinguish between original human argumentation and typical AI argumentation.
Table 2 shows the dataset statistics. Human essays were collected from student submissions. AI essays were those created following AIDE’s generation protocol. Specifically, 3-shot prompting was used, with three randomly selected human essays on the same topic as in-context examples, generating essays based on the same assignment instructions and reference materials. After generation, the following post-processing was applied: i) deduplication with a threshold of 85% cosine similarity using Universal Sentence Encoder embeddings, ii) manual removal of outputs that students would obviously not submit, iii) normalization of formatting and special characters, and iv) statistical matching of word count, sentence length, and function word distributions to human essays.
@@
For the five-element extraction and quality evaluation, this study used GPT-4.1-mini [22]. GPT-4.1-mini has high instruction-following capability while having low API cost, making it suitable for processing large-scale datasets. Other parameters follow the settings described in Section 3.

### v3

@@ 3.1 Framework Overview
AROA is a structure-oriented originality assessment framework that evaluates essays based on the originality of thinking rather than surface fluency or author attribution. Fig. 1 shows the overall processing pipeline of AROA. The design philosophy of AROA is based on Toulmin’s argumentation model [Toulmin, 2003] and integrates insights from critical thinking education [Ennis, 1985, Facione, 1990], deep learning theory [Biggs, 1986], and knowledge construction theory [Scardamalia and Bereiter, 2006].
@@
These four perspectives capture different dimensions of originality: how arguments are made (structure), what is argued (claims and evidence), and how deeply arguments are developed (cognitive depth). The following sections detail each processing stage.
@@ 3.2 Argument Structure Extraction
For each essay $x$, we define a structural representation consisting of the following five elements:
@@
The definition of each element is shown in Table 1. This five-element representation adapts Toulmin’s model [Toulmin, 2003] by integrating Warrant and Backing into Reasons and omitting Qualifier, improving practicality in educational settings and consistency of LLM extraction.
@@
These elements form hierarchical relationships: Reasons support Claims, Evidence backs up Reasons, Counterarguments oppose Claims, and Rebuttals respond to Counterarguments. This structure captures not only element presence but also the logical composition of argumentation.
An LLM is used for structure extraction with a system prompt specifying the task definition and guidelines:
@@
These guidelines ensure faithful extraction, unified element granularity, and handling of missing elements. The essay text is provided in the user prompt with instructions for JSON-formatted output including element lists and relational information. The temperature parameter was set to $T=0$ for deterministic output.
From the extracted structure, two derived features are calculated. The depth of reasoning $\text{Depth}(x)$ represents hierarchical complexity:
@@
This definition distinguishes simple argumentation (depth 1) from complex argumentation with evidence or counterargument-rebuttal structure (depth 3). The number of counterargument-rebuttal matched pairs $n_{\text{matched}}$ is counted from responds_to relationships, or estimated as $n_{\text{matched}}=\min(n_{A},n_{B})$ when relationships are not explicitly extracted.
@@
This study used a pre-trained Sentence-BERT model [Reimers and Gurevych, 2019] to obtain $d=768$-dimensional embeddings. These representations serve as the basis for semantic rarity assessment in subsequent sections.
@@ 3.3 Structural Rarity Assessment
The structural rarity score $S_{\text{struct}}$ quantifies how rare an essay’s argument structure is within the corpus. A 12-dimensional structural feature vector $\boldsymbol{\phi}(x)$ is constructed from each essay $x$:
@@
Here, $n_{C},n_{R},n_{E},n_{A},n_{B}$ are the counts of each element, $n_{R}/\max(n_{C},1)$ is the number of reasons per claim, and $n_{E}/\max(n_{R},1)$ is the number of evidences per reason. $\text{Depth}(x)$ and $n_{\text{matched}}$ are the depth of reasoning and the number of counterargument-rebuttal matched pairs defined in the previous section. $\text{Branches}(x)=\max(n_{C},1)\times\max(n_{R},1)$ represents the number of argument branches, capturing multifaceted argument structures that support multiple claims with multiple reasons. $\mathbb{I}_{n_{A}>0}$ and $\mathbb{I}_{n_{B}>0}$ are indicator functions showing the presence or absence of counterarguments and rebuttals. These features capture quantitative (element counts), structural (ratios, depth, branches), and critical thinking aspects (counterargument-rebuttal correspondence).
The structural rarity score is defined as the negative log-density in the feature vector space, corresponding to self-information in information theory. Prior to density estimation, standardization is performed on the set of feature vectors across the entire corpus $\Phi=\{\boldsymbol{\phi}(x_{i})\}_{i=1}^{N}$:
@@
Here, $\boldsymbol{\mu}_{\Phi}$ and $\boldsymbol{\sigma}_{\Phi}$ are the corpus-wide mean and standard deviation of each feature. Dimensionality is then reduced to 3 using PCA to mitigate the curse of dimensionality, and Kernel Density Estimation (KDE) with Scott’s rule bandwidth is applied:
@@
Here, $\hat{p}(\boldsymbol{\phi}^{\prime}(x))$ is the KDE density estimate after standardization and PCA. Essays with typical structures receive low scores, while those with rare structures receive high scores.
@@ 3.4 Semantic Rarity Assessment
The semantic rarity scores $S_{\text{claim}}$ and $S_{\text{evid}}$ quantify how rare the content of claims and evidence are within the corpus. While structural rarity evaluates how arguments are made, semantic rarity evaluates what is argued.
@@ 3.5 Cognitive Depth Assessment
The cognitive depth score $S_{\text{move}}$ evaluates the depth of reasoning and completeness of response to counterarguments. While rarity assessments measure whether something is rare, cognitive depth evaluates how deeply arguments are developed.
@@
The first component, $n_{\text{matched}}$, is the number of counterargument-rebuttal matched pairs as defined in Section 3.2. The second component, $\text{Depth}(x)$, is the depth of reasoning as defined in Section 3.2. The third component, $\text{Coverage}(x)$, is the counterargument coverage rate, representing the proportion of presented counterarguments that are responded to with rebuttals:
@@
This metric captures whether counterarguments are not only raised but also appropriately responded to. This study used $\alpha=\beta=\gamma=1.0$, treating the three components equally. These weights can be adjusted according to educational objectives.
@@ 3.6 Quality Adjustment
The quality score $Q(x)$ prevents essays that are original but logically flawed from receiving high scores. The quality score consists of two components: Coherence and Logical Validity:
@@
This framework intentionally excludes fluency from quality evaluation, as LLM-generated text generally has high fluency, and including it would favor AI-generated text. Each component is evaluated using an LLM. For coherence evaluation, judgment is made based on three criteria: i) whether there are contradictions between claims and reasons, ii) whether there are contradictions between reasons and evidence, and iii) whether there are contradictions among multiple claims. Output is in three levels: 1.0 (no contradictions), 0.5 (partial contradictions), and 0.0 (serious contradictions). For logical validity evaluation, judgment is made based on three criteria: i) whether reasons logically support claims, ii) whether evidence appropriately backs up reasons, and iii) whether there are logical leaps or fallacies. Output is in three levels: 1.0 (valid), 0.5 (partial problems), and 0.0 (serious defects).
In both evaluations, the prompt explicitly instructs not to evaluate fluency or style, focusing evaluation on logical structure only.
@@ 3.7 Final Score Computation
The four sub-scores ($S_{\text{struct}}$, $S_{\text{claim}}$, $S_{\text{evid}}$, $S_{\text{move}}$) have different scales, so each is z-score normalized within the corpus:
@@
Here, $\mu_{\text{comp}}$ and $\sigma_{\text{comp}}$ are the corpus-wide mean and standard deviation of each component.
@@
Here, $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}$ are weights satisfying $\sum_{i=1}^{4}\lambda_{i}=1$. This study used equal weights $\lambda_{i}=0.25$.
@@
This multiplicative design ensures that logically flawed essays ($Q(x)\approx 0$) receive low final scores regardless of originality, positioning quality as a prerequisite for originality. Algorithm 1 summarizes the overall processing procedure of AROA.
@@
17:  if $n_{A}(x_{i})>0$ then
18:   $\text{Coverage}(x_{i})\leftarrow n_{\text{matched}}(x_{i})/n_{A}(x_{i})$
19:  else
20:   $\text{Coverage}(x_{i})\leftarrow 0$
21:  end if
22:  $S_{\text{move}}(x_{i})\leftarrow\alpha\cdot n_{\text{matched}}(x_{i})+\beta\cdot\text{Depth}(x_{i})+\gamma\cdot\text{Coverage}(x_{i})$
23:end for
24:// Stage 4: Quality Adjustment
25:for each $x_{i}\in\mathcal{D}$ do
26:  $Q_{\text{coherence}}(x_{i})\leftarrow\text{LLM\_Evaluate}(x_{i},\text{``coherence''})$
27:  $Q_{\text{logical}}(x_{i})\leftarrow\text{LLM\_Evaluate}(x_{i},\text{``logical''})$
28:  $Q(x_{i})\leftarrow w_{\text{coh}}\cdot Q_{\text{coherence}}(x_{i})+w_{\text{log}}\cdot Q_{\text{logical}}(x_{i})$
29:end for
30:// Stage 5: Final Score Computation
31:for each component $\text{comp}\in\{\text{struct},\text{claim},\text{evid},\text{move}\}$ do
32:  $\mu_{\text{comp}},\sigma_{\text{comp}}\leftarrow\text{Mean}(\{S_{\text{comp}}(x_{i})\}),\text{Std}(\{S_{\text{comp}}(x_{i})\})$
33:  $\tilde{S}_{\text{comp}}(x_{i})\leftarrow(S_{\text{comp}}(x_{i})-\mu_{\text{comp}})/\sigma_{\text{comp}}$ for all $i$
@@
35:for each $x_{i}\in\mathcal{D}$ do
36:  $\tilde{S}_{\text{orig}}(x_{i})\leftarrow\sum_{j=1}^{4}\lambda_{j}\tilde{S}_{j}(x_{i})$
37:  $S_{\text{final}}(x_{i})\leftarrow\tilde{S}_{\text{orig}}(x_{i})\times Q(x_{i})$
38:end for
39:return $\{S_{\text{final}}(x_{i})\}_{i=1}^{N}$
@@ 4.1 Dataset
This study used the AI Detection for Essays Dataset (AIDE) [Burleigh, 2024], a publicly available dataset on Kaggle constructed by Vanderbilt University and The Learning Agency Lab for AI-generated text detection in educational settings. AIDE contains argumentative essays written by middle and high school students alongside essays generated by multiple LLMs. The essays address two topics—Car-free cities and Electoral College—both requiring students to read given source materials and argue their own position. This dataset was selected because it contains numerous human essays on the same topic, suitable for rarity-based evaluation, and includes both human and AI essays for comparative analysis.
Table 2 shows the dataset statistics. Human essays were collected from student submissions. The original AIDE dataset contains AI essays generated using multiple LLMs including PaLM 2 (text-bison, text-unicorn), Gemini 1.0 (Pro, Ultra), and GPT-4, with 3-shot prompting using randomly selected human essays as in-context examples and random sampling at temperature 1.0–1.6. Post-processing included deduplication (85% cosine similarity threshold), removal of implausible outputs, formatting normalization, and statistical matching of text characteristics to human essays; approximately 10% of originally generated essays were retained. Since the original AIDE dataset contained fewer AI essays than needed for balanced comparison, we supplemented the AI essay collection to reach 1,000 essays (500 per topic) by generating additional essays using GPT-4.1-mini following the same generation protocol as AIDE: 3-shot prompting with randomly selected human essays as in-context examples, the same assignment instructions and source materials, and equivalent post-processing steps.
@@
For the argument structure extraction and quality evaluation in the AROA pipeline, this study used GPT-4.1-mini [OpenAI, 2024], which offers high instruction-following capability at low API cost, making it suitable for processing large-scale datasets. Other parameters follow the settings described in Section 3.